# Is Ego Status All You Need for Open-Loop End-to-End Autonomous Driving?

### v1

@@ Abstract
End-to-end autonomous driving recently emerged as a promising research direction to target autonomy from a full-stack perspective. Along this line, many of the latest works follow an open-loop evaluation setting on nuScenes to study the planning behavior. In this paper, we delve deeper into the problem by conducting thorough analyses and demystifying more devils in the details. We initially observed that the nuScenes dataset, characterized by relatively simple driving scenarios, leads to an under-utilization of perception information in end-to-end models incorporating ego status, such as the ego vehicle’s velocity. These models tend to rely predominantly on the ego vehicle’s status for future path planning. Beyond the limitations of the dataset, we also note that current metrics do not comprehensively assess the planning quality, leading to potentially biased conclusions drawn from existing benchmarks. To address this issue, we introduce a new metric to evaluate whether the predicted trajectories adhere to the road. We further propose a simple baseline able to achieve competitive results without relying on perception annotations. Given the current limitations on the benchmark and metrics, we suggest the community reassess relevant prevailing research and be cautious whether the continued pursuit of state-of-the-art would yield convincing and universal conclusions. Code and models are available at https://github.com/NVlabs/BEV-Planner
* Work done during an internship at NVIDIA.
@@ 3 Proposed BEV-Planner
In fact, ST-P3, a previous method that often serves as a baseline, uses partially incorrect GT data during training and evaluation [^2] https://github.com/OpenDriveLab/ST-P3/issues/24. Consequently, when conducting comparisons between other methods and ST-P3, the validity of the conclusions drawn must be carefully evaluated. Therefore, in this paper, it is necessary for us to redesign a baseline method to compare with existing methods. At the same time, to better explore the impact of ego status, we also need a relatively clear baseline method. Based on these considerations, we have designed a very simple baseline in this paper, named BEV-Planner, as shown in the  Fig. 1(c). For our pipeline, we first generate the BEV feature and concatenate it with history BEV features, mainly following the previous method [12, 14, 21]. Please note while concatenating BEV features from different timesteps, we didn’t perform feature alignment. After obtaining the BEV features, we directly perform a cross-attention [36] between the BEV features and the ego query, which is a learnable embedding. The final trajectories are predicted based on the refined ego query through MLPs. The process can be formulated as follows:
@@ 4.2 Metrics
Collision Rate.
$CR(t)=\frac{\sum_{i=0}^{N}\mathbb{I}_{i}}{N},N=t/0.5,$ (2)
$CR(t)={(\sum_{i=0}^{N}\mathbb{I}_{i}})>0,N=t/0.5.$ (3)
Intersection Rate with Road Boundary.
In this study, to more comprehensively assess the quality of predicted trajectories, we employed a new metric that calculates the intersection rate between the predicted trajectories and road boundaries. Staying on the road is vital for the safety of autonomous driving systems, yet existing evaluation metrics overlook the integration of map priors. Intuitively, safe trajectories should avoid interactions with road boundaries. Interactions with road boundaries typically indicate the possibility of leaving the drivable area, which can pose safety hazards. We recognize that certain annotated road boundaries on nuScenes are indeed traversable, and ground truth trajectories may intersect with these boundaries under specific conditions. However, from a statistical viewpoint, this metric can effectively represent the overall rationality of the model’s predicted trajectories. The implementation of the intersection rate with road boundary metric is informed by the collision rate. To facilitate this, we rasterize the road boundary using a resolution of 0.1 meters.
@@
Considering the absence of an official implementation of evaluation metrics, existing methods may vary in the details of their metric implementations. In this paper, we used the official open-source repositories of other methods to generate trajectory prediction results and then applied the same metric implementation to evaluate all methods, ensuring fairness. Addressing concerns raised by AD-MLP [45] regarding the potential for false collisions due to a coarse-grained grid size (0.5m), we adopt a finer default grid size of 0.1m in our work to mitigate this issue.
ID Method Ego Status L2 (m) $\downarrow$ Collision (%) $\downarrow$ Intersection (%) $\downarrow$ ckpt. source
@@
Table 2: The VAD-base model’s robustness to images and ego status. To ascertain the impact of perceptual information and ego status on the ultimate planning performance, we systematically introduced noise into each component separately. We utilize the official VAD-Base checkpoint that uses ego status in its planner module. *: the results of VAD-Base without ego status in its planner. We can observe that introducing corruption to images markedly affects the perception results, especially in the case of using blank images; nonetheless, this does not markedly disrupt the ultimate planning results. In contrast to the minor impact of image corruption on planning, modifications to the ego vehicle’s velocity have a significant effect on the planning results. Experimental results reveal that in an end-to-end model incorporating both ego status and perceptual information, decision-making is disproportionately influenced by ego status, thereby substantially increasing the model’s safety risks. : The collision rate is not precise as the ego car may have departed from the local BEV area. When the input velocity is zero, the model produces almost stationary trajectories, resulting in excellent performance in the intersection with road boundary metrics. This can be seen as a limitation of the intersection metric.
Figure 4: We exhibit the predicted trajectories of the VAD model (incorporating ego status in its planner) under various image corruptions. All trajectories within a given scene (spanning 20 seconds) are presented in the global coordinate system. Each triangular marker signifies a ground truth trajectory point of the ego vehicle, with different colors representing distinct timesteps. Notably, the model’s predicted trajectory maintains plausibility, even when blank images serve as input. The trajectories within the red boxes, however, are suboptimal, as further elucidated in  Fig. 5. While corruptions were applied to all surround-view images, for the sake of visualization, only the corresponding front-view images at the initial timestep are displayed.
@@ 4.3 Discussion
While focusing solely on previous metrics L2 distance and collision rate, it is observed that the simple strategy (ID-7), which simply continues straight at the current velocity, achieves surprisingly good results. The Ego-MLP model, which didn’t leverage perception clues, actually is on par with UniAD and VAD, which use a more complex pipeline. From another perspective, it is observed that existing methods can only match the performance of Ego-MLP when the ego vehicle’s status is incorporated into the planner. In contrast, reliance solely on camera inputs leads to results significantly inferior to those achieved by Ego-MLP. Considering these observations, we may tentatively infer an intriguing conclusion: utilizing a combination of sensory information and ego status appears to yield results comparable to those achieved by employing ego status alone. Therefore, in models that integrate both ego vehicle status and perception information, a pertinent question arises: What specific role does the perception information, acquired from camera inputs, play within the final planning module?
Ego Status vs.  Perceptual Information
Undoubtedly, perceptual information constitutes the indispensable foundation of all autonomous driving systems, with ego status additionally offering crucial data such as the vehicle’s velocity and acceleration to aid the system’s decision-making process. Incorporating both perceptual information and ego status for the ultimate planning should indeed be a judicious strategy within an end-to-end autonomous driving system. However, as shown in Tab. 1, relying solely on ego status can yield planning results that are on par with or even superior to those methods that utilize both ego status and perception modules on previous L2 or collation rate metrics. To ascertain the roles that perceptual information and ego status play in the final planning process, we introduced varying degrees of perturbation to the images and the ego status, as shown in the Tab. 2. We use the official VAD model (which leverages ego status in the planner module) as the base model. It is observable that when disturbances are added to the images, the results of planning marginally decrease and may even exhibit improvement, while perceptual performance significantly deteriorates. Surprisingly, even when blank images are used as input, leading to the complete breakdown of the perception module, the model’s planning capabilities remain largely unaffected. The corresponding visualization results are as illustrated in the  Fig. 4. In contrast to the model’s remarkable robustness to variations in image inputs, it exhibits considerable sensitivity to ego status. Upon altering the velocity of the ego car, we can observe that the planning results of the model are significantly impacted, as shown in Fig. 6. Setting the ego car’s speed to 100 m/s results in the model generating wildly impractical planning trajectories. We posit that an autonomous driving system displaying such heightened sensitivity to ego status information harbors considerable safety risks. Furthermore, with planning results being predominantly dictated by ego status, the functions of other modules in the model cannot be reflected. For example, while comparing VAD(ID-6) and BEV-Planner++ (ID-12), we can observe that they obtain basically similar results in terms of L2 and collision rate. Is it justifiable to assert that our BEV-Planner++ design, characterized by its simplicity and effectiveness, can attain comparable outcomes to other more intricate methodologies, even in the absence of utilizing perception data? In fact, as the performance of the final planning module is predominantly influenced by the ego vehicle status, the design of other components does not significantly affect the planning outcomes. Consequently, we argue that methods utilizing ego status are not directly comparable and conclusions should not be drawn from such comparisons.
How about not using ego status?
Neglected Ego Status in Perception Stage.
[…]
BEVFormer 41.6 51.7
BEVFormer 41.3 51.5
Method Ego Status L2 (m) $\downarrow$ Collision (%) $\downarrow$ Intersection (%) $\downarrow$
@@
Table 4: While adding map perception task into the BEV-Planner method, we can observe that the model obtains worse L2 distance and collision rate performance, but achieves better intersection rate with road boundaries. We use a pretrained map perception checkpoint as the initialization of the BEV-Planner (init*) and BEV-Planner+Map model.
@@
People might wonder why our BEV-Planner, without using additional perception tasks (including Depth, HD map, Tracking, etc.) and ego status, achieves better results in L2 distance and collision rate than other methods (ID-1 and 4). Since our BEV-Planner performs poorly in terms of intersection with road boundaries, what would happen if we added map perception tasks to our baseline? To address these questions, we designed a “BEV-Planner+Map” model by introducing a map perception task into our pipeline, mainly following the designs of UniAD. As shown in Tab. 4, when map perception is introduced, the model exhibits poorer results in terms of L2 distance and collision rate metrics. The only aspect that aligns with our expectations is that the introduction of map perception significantly reduces the intersection rate with road boundaries. Through a comparison of BEV-Planner with BEV-Planner (init*), we observe that the use of map-pretrained weights can enhance performance. This finding implies that the decrease in L2 and Collision rate observed with the integration of MapFormer in “BEV-Planner+Map” is not due to the pretrained weights. We posit that in most straight-driving scenarios, the addition of lane information may not yield markedly effective information and could indeed introduce some degree of interference. To verify our hypothesis, we evaluated the performance of these methods under varying driving commands. As shown in Tab. 5 and Tab. 6, adding map information significantly increases the L2 distance error and collision rate with going straight commands. In contrast, for turning scenarios, the incorporation of map information effectively reduces the collision rate. Based on the above observations, we can tentatively draw the following conclusions:
@@
Table 7: P.P. indicates the post-processing optimization module of UniAD. We use the officially released weights of UniAD for this ablation study. By default, UniAD incorporates a post-processing step to refine the predicted trajectory from the end-to-end model, aiming to mitigate collisions with other agents. Nonetheless, this approach is limited by this singular optimization objective, which overlooks additional safety-critical factors in autonomous driving, such as lane adherence. Our new metrics, the intersection rate with road boundary reveal that UniAD’s post-processing substantially increases the risk of the ego vehicle running off the road.
@@
The preceding methodology primarily centered around the L2 distance and collision rate metrics. Our discussion thus far has been largely concentrated on these two metrics. What we want to emphasize is that these two metrics, L2 distance and collision rate, only reflect a partial aspect of a model’s planning capabilities. It is not advisable to assess the quality of a model based exclusively on these two metrics. In this paper, we introduce a new metric to evaluate the model’s comprehension and adherence to the map: Intersection Rate with Road Boundary. As shown in Tab. 1, we can observe that the GoStright strategy frequently intersects with road boundaries, which is in line with our expectations. In terms of this new metric, Ego-MLP performs worse than UniAD and VAD. Our method BEV-Planner, performs the worst on this metric because it does not use any map information. We assert that the evaluation of open-loop autonomous driving systems necessitates a more diverse and stringent evaluation framework. This would enable a more accurate reflection of these systems’ capabilities and limitations.
@@
Failure Cases
Although the majority of scenarios in the nuScenes dataset are relatively straightforward, it does include certain challenging scenes, notably those involving continuous cornering. As shown in Fig. 10, we can observe that methods with various settings all yielded suboptimal predicted trajectories when navigating high-curvature bends. For challenging scenarios like cornering, where the system must continuously make evolving decisions, evaluating open-loop autonomous driving systems poses a significant challenge. One limitation of open-loop methods is that they do not suffer from cumulative errors. In detail, in the case of an extremely erroneous trajectory predicted at a given timestep, the trajectory starting point for the next timestep is still based on the GT trajectory. The metric we utilize, intersection with road boundary, is adept at identifying low-quality trajectories. However, an appropriate metric that can effectively highlight high-quality trajectories remains an intriguing direction for further exploration.
@@ 5 Conclusion
There are trade-offs between different planning metrics. Designing an integrated evaluation system for open-loop evaluation presents a significant challenge. Although our baseline method excels in terms of L2 distance and collision rate, its performance is not exceptional in the intersection rate with road boundary metric, primarily because our approach does not utilize any perception annotations, such as HD maps.

### v2

@@ Abstract
End-to-end autonomous driving recently emerged as a promising research direction to target autonomy from a full-stack perspective. Along this line, many of the latest works follow an open-loop evaluation setting on nuScenes to study the planning behavior. In this paper, we delve deeper into the problem by conducting thorough analyses and demystifying more devils in the details. We initially observed that the nuScenes dataset, characterized by relatively simple driving scenarios, leads to an under-utilization of perception information in end-to-end models incorporating ego status, such as the ego vehicle’s velocity. These models tend to rely predominantly on the ego vehicle’s status for future path planning. Beyond the limitations of the dataset, we also note that current metrics do not comprehensively assess the planning quality, leading to potentially biased conclusions drawn from existing benchmarks. To address this issue, we introduce a new metric to evaluate whether the predicted trajectories adhere to the road. We further propose a simple baseline able to achieve competitive results without relying on perception annotations. Given the current limitations on the benchmark and metrics, we suggest the community reassess relevant prevailing research and be cautious about whether the continued pursuit of state-of-the-art would yield convincing and universal conclusions. Code and models are available at https://github.com/NVlabs/BEV-Planner.
@@ 3 Proposed BEV-Planner
In fact, ST-P3 [12], a previous method that often serves as a baseline, uses partially incorrect GT data during training and evaluation [^2] https://github.com/OpenDriveLab/ST-P3/issues/24. Consequently, when conducting comparisons between other methods and ST-P3, the validity of the conclusions drawn must be carefully evaluated. Therefore, in this paper, it is necessary for us to redesign a baseline method to compare with existing methods. At the same time, to better explore the impact of ego status, we also need a relatively clear baseline method. Based on these considerations, we have designed a very simple baseline in this paper, named BEV-Planner, as shown in the  Fig. 1(c). For our pipeline, we first generate the BEV feature and concatenate it with history BEV features, mainly following the previous method [12, 14, 21]. Please note while concatenating BEV features from different timesteps, we didn’t perform feature alignment. After obtaining the BEV features, we directly perform a cross-attention [36] between the BEV features and the ego query, which is a learnable embedding. The final trajectories are predicted based on the refined ego query through MLPs. The process can be formulated as follows:
@@ 4.2 Metrics
In the Appendix, we will introduce the shortcomings of the currently commonly used collision rate and another metric used to evaluate the smoothness of predicted trajectory.
Curb Collision Rate (CCR).
In this study, to more comprehensively assess the quality of predicted trajectories, we employed a new metric that calculates the collision rate between the predicted trajectories and curbs (road boundaries). Staying on the road is vital for the safety of autonomous driving systems, yet existing evaluation metrics overlook the integration of map priors. Intuitively, safe trajectories should avoid collision with curbs. Our Curb Collision Rate (CCR) typically indicate the possibility of leaving the drivable area, which can pose safety hazards. We recognize that certain annotated road boundaries on nuScenes are indeed traversable, and ground truth trajectories may intersect with these boundaries under specific conditions. However, from a statistical viewpoint, this metric can effectively represent the overall rationality of the model’s predicted trajectories. The implementation of the CCR metric is informed by the collision rate. To facilitate this, we rasterize the road boundary using a resolution of 0.1 meters. More details are in the Appendix.
@@
Given the lack of a standardized approach to assessment metrics, methodologies might differ in the nuances of their metric executions. In this study, we utilized the official open-source repositories from various methodologies to produce predicted trajectories. Subsequently, we employed a consistent metric implementation across all methods for evaluation, guaranteeing equity. Addressing concerns raised by AD-MLP [45] regarding the potential for false collisions due to a coarse-grained grid size (0.5m), we adopt a finer default grid size of 0.1m in our work to mitigate this issue.
ID Method Ego Status L2 (m) $\downarrow$ Collision (%) $\downarrow$ CCR (%) $\downarrow$ ckpt. source
@@
Method Img Ego Status L2 (m) $\downarrow$ Collision (%) $\downarrow$ CCR (%) $\downarrow$ Det. Map
@@
Table 2: The VAD-base model’s robustness to images and ego status. To ascertain the impact of perceptual information and ego status on the ultimate planning performance, we systematically introduced noise into each component separately. We utilize the official VAD-Base checkpoint that uses ego status in its planner module. *: the results of VAD-Base without ego status in its planner. We can observe that introducing corruption to images markedly affects the perception results, especially in the case of using blank images; nonetheless, this does not markedly disrupt the ultimate planning results. In contrast to the minor impact of image corruption on planning, modifications to the ego vehicle’s velocity have a significant effect on the planning results. Experimental results reveal that in an end-to-end model incorporating both ego status and perceptual information, decision-making is disproportionately influenced by ego status, thereby substantially increasing the model’s safety risks. : The collision rate is not precise as the ego car may have departed from the local BEV area. When the input velocity is zero, the model produces almost stationary trajectories, resulting in excellent performance in the CCR. This can be seen as a limitation of the CCR metric.
Figure 3: We exhibit the predicted trajectories of the VAD model (incorporating ego status in its planner) under various image corruptions. All trajectories within a given scene (spanning 20 seconds) are presented in the global coordinate system. Each triangular marker signifies a ground truth trajectory point of the ego vehicle, with different colors representing distinct timesteps. Notably, the model’s predicted trajectory maintains plausibility, even when blank images serve as input. The trajectories within the red boxes, however, are suboptimal, as further elucidated in Appendix. While corruptions were applied to all surround-view images, for the sake of visualization, only the corresponding front-view images at the initial timestep are displayed.
@@ 4.3 Discussion
While focusing solely on previous metrics L2 distance and collision rate, it is observed that the simple strategy (ID-7), which simply continues straight at the current velocity, achieves surprisingly good results. The Ego-MLP model, which didn’t leverage perception clues, is actually on par with UniAD and VAD, which use more complex pipelines. From another perspective, it is observed that existing methods can only match the performance of Ego-MLP when the ego vehicle’s status is incorporated into the planner. In contrast, reliance solely on camera inputs leads to results significantly inferior to those achieved by Ego-MLP. Considering these observations, we may tentatively infer an intriguing conclusion: utilizing a combination of sensory information and ego status appears to yield results comparable to those achieved by employing ego status alone. Therefore, in models that integrate both ego vehicle status and perception information, a pertinent question arises: What specific role does the perception information, acquired from camera inputs, play within the final planning module?
Method Ego Status L2 (m) $\downarrow$ Collision (%) $\downarrow$ CCR (%) $\downarrow$
@@
Table 3: While adding map perception task into the BEV-Planner method, we can observe that the model obtains worse L2 distance and collision rate performance, but achieves better CCR. We use a pretrained map perception checkpoint as the initialization of the BEV-Planner (init*) and BEV-Planner+Map model.
@@
Ego Status vs.  Perceptual Information
Undoubtedly, perceptual information constitutes the indispensable foundation of all autonomous driving systems, with ego status additionally offering crucial data such as the vehicle’s velocity and acceleration to aid the system’s decision-making process. Incorporating both perceptual information and ego status for the ultimate planning should indeed be a judicious strategy within an end-to-end autonomous driving system. However, as shown in Tab. 1, relying solely on ego status can yield planning results that are on par with or even superior to those methods that utilize both ego status and perception modules on previous L2 or collation rate metrics. To ascertain the roles that perceptual information and ego status play in the final planning process, we introduced varying degrees of perturbation to the images and the ego status, as shown in the Tab. 2. We use the official VAD model (which leverages ego status in the planner module) as the base model. It is observable that when disturbances are added to the images, the results of planning marginally decrease and may even exhibit improvement, while perceptual performance significantly deteriorates. Surprisingly, even when blank images are used as input, leading to the complete breakdown of the perception module, the model’s planning capabilities remain largely unaffected. The corresponding visualization results are as illustrated in the  Fig. 3. In contrast to the model’s remarkable robustness to variations in image inputs, it exhibits considerable sensitivity to ego status. Upon altering the velocity of the ego car, we can observe that the planning results of the model are significantly impacted, as shown in Fig. 4. Setting the ego car’s speed to 100 m/s results in the model generating wildly impractical planning trajectories. We posit that an autonomous driving system displaying such heightened sensitivity to ego status information harbors considerable safety risks. Furthermore, with planning results being predominantly dictated by ego status, the functions of other modules in the model cannot be reflected. For example, while comparing VAD(ID-6) and BEV-Planner++ (ID-12), we can observe that they obtain basically similar results in terms of L2 and collision rate. Is it justifiable to assert that our BEV-Planner++ design, characterized by its simplicity and effectiveness, can attain comparable outcomes to other more intricate methodologies, even in the absence of utilizing perception data? In fact, as the performance of the final planning module is predominantly influenced by the ego vehicle status, the design of other components does not significantly affect the planning results. Consequently, we argue that methods utilizing ego status are not directly comparable and conclusions should not be drawn from such comparisons.
How about not using ego status?
Neglected Ego Status in Perception Stage.
In fact, existing methods [16, 43] ignore the impact of using ego status on planning in BEV Encoder. More details are in the Appendix.
@@
People might wonder why our BEV-Planner, without using additional perception tasks (including Depth, HD map, Tracking, etc.) and ego status, achieves better results in L2 distance and collision rate than other methods (ID-1 and 4). Since our BEV-Planner performs poorly in terms of CCR, what would happen if we added map perception tasks to our baseline? To address these questions, we designed a “BEV-Planner+Map" model by introducing a map perception task into our pipeline, mainly following the designs of UniAD. As shown in Tab. 3, when map perception is introduced, the model exhibits poorer results in terms of L2 distance and collision rate metrics. The only aspect that aligns with our expectations is that the introduction of map perception significantly reduces the CCR. Through a comparison of BEV-Planner with BEV-Planner (init*), we observe that the use of map-pretrained weights can enhance performance. This finding implies that the decrease in L2 and Collision rate observed with the integration of MapFormer in “BEV-Planner+Map" is not due to the pretrained weights. We posit that in most straight-driving scenarios, the addition of lane information may not yield markedly effective information and could indeed introduce some degree of interference. To verify our hypothesis, we evaluated the performance of these methods under varying driving commands. As shown in Tab. 4 and Tab. 5, adding map information significantly increases the L2 distance error and collision rate with going straight commands. In contrast, for turning scenarios, the incorporation of map information effectively reduces the collision rate. Based on the above observations, we can tentatively draw the following conclusions:
@@
Table 6: P.P. indicates the post-processing optimization module of UniAD. We use the officially released weights of UniAD for this ablation study. By default, UniAD incorporates a post-processing step to refine the predicted trajectory from the end-to-end model, aiming to mitigate collisions with other agents. Nonetheless, this approach is limited by this singular optimization objective, which overlooks additional safety-critical factors in autonomous driving, such as lane adherence. Our new metrics, the CCR reveal that UniAD’s post-processing substantially increases the risk of the ego vehicle running off the road.
@@
The preceding methodology primarily centered around the L2 distance and collision rate metrics. Our discussion thus far has been largely concentrated on these two metrics. What we want to emphasize is that these two metrics, L2 distance and collision rate, only reflect a partial aspect of a model’s planning capabilities. It is not advisable to assess the quality of a model based exclusively on these two metrics. In this paper, we introduce a new metric to evaluate the model’s comprehension and adherence to the map: Curb Collision Rate (CCR). As shown in Tab. 1, we can observe that the GoStright strategy frequently intersects with road boundaries, which is in line with our expectations. In terms of this new metric, Ego-MLP performs worse than UniAD and VAD. Our method BEV-Planner, performs the worst on this metric because it does not use any map information. This suggests that relying on past metrics to judge the superiority of different open-loop methods is biased.
@@
What we wish to underscore is that the primary intention behind proposing CCR is to illuminate the inadequacies within the existing evaluation systems. However, even with the incorporation of CCR, open-loop evaluation systems still encounter numerous challenges. We assert that the evaluation of open-loop autonomous driving systems necessitates a more diverse and stringent evaluation framework. This would enable a more accurate reflection of these systems’ capabilities and limitations.
@@ 5 Conclusion
There are trade-offs between different planning metrics. Designing an integrated evaluation system for open-loop evaluation presents a significant challenge. Although our baseline method excels in terms of L2 distance and collision rate, its performance is not exceptional in the CCR, primarily because our approach does not utilize any perception annotations, such as HD maps.
Acknowledgments
Thanks to Bencheng Liao and Shaoyu Chen for providing helpful information and the model weights of VAD. Tong Lu and Zhiqi Li are supported by the National Natural Science Foundation of China (Grant No. 62372223) and China Mobile Zijin Innovation Insititute (No. NR2310J7M). Zhiqi Li is also supported by the NVIDIA Graduate Fellowship Program.